OpenStudy (anonymous):

Probability: Nicole is taking a 5-question multiple choice test. If she guesses every answer, and each question has 4 choices, what is the probability that she gets at most one wrong answer?

OpenStudy (amistre64):

P(0wrong) + P(1wrong) is part of it ..... kinda fuzzy how to do the rest wwwr; success= 1/4 , fail = 3/4 per question , hnn

OpenStudy (amistre64):

the probability of getting a question wrong seems to me to be: 1/5 * 3/4 = 3/20

OpenStudy (amistre64):

wwwww wwwwr wwwrw wwrww wrwww rwwww wwwrr wwrwr wrwwr 1,5,10,10,5,1 = 32 possible outcomes rrrrr = 1 possibility rrrrw = 5 possibilities so maybe 6/32 has a role in there

OpenStudy (amistre64):

can you refresh my memory any, how did you try to work it?

OpenStudy (anonymous):

i used the formula nCr(p)^r(q)^n-r

OpenStudy (amistre64):

n=20 right? r = 0 or 1

OpenStudy (anonymous):

im not sure ): i put 5 for n and yea i got 0 and 1 for r

OpenStudy (amistre64):

there are 20 answers in all to choose from; you want the probability of getting 0 wrong + 1 wrong

OpenStudy (amistre64):

lets try that 1(3/4)^0 (1/4)^(20) + 20(3/4)^1 (1/4)^(19) it aint 20, your 5 is right 1(3/4)^0 (1/4)^(5) + 20(3/4)^1 (1/4)^(4) http://www.wolframalpha.com/input/?i=1%283%2F4%29%5E0+%281%2F4%29%5E%285%29+%2B+5%283%2F4%29%5E1+%281%2F4%29%5E%284%29

OpenStudy (amistre64):

errr, i forgot to change the 20 to a 5 in the post :)

OpenStudy (anonymous):

hold on ill try it o:

OpenStudy (amistre64):

P(0wrong) = 5C0 w^0 r^5 P(1wrong) = 5C1 w^1 r^4 P(at most 1 wrong) = P(0) + P(1)

OpenStudy (anonymous):

We can reason through this: P(0 wrong) = P(all right) = (1/4)^5 P(1 wrong) = (3/4)(1/4)^4 but this can happen in 5 different ways So out solution is (1/4)^5+5(3/4(1/4)^4))

OpenStudy (anonymous):

thats what i put in but i dont get the decimal im supposed to ):

OpenStudy (anonymous):

(1/4)^5+5(3/4(1/4)^4)) = 1/64 = 0.015625

OpenStudy (anonymous):

wait i think i mixed up p and q ! what were they?

OpenStudy (anonymous):

Sorry I'm not good at combinatorics so I can't tell you how to do it formally. I just reasoned it out using basic probability

OpenStudy (anonymous):

oh its alright, thanks so much tho!!

OpenStudy (anonymous):

Glad to help